OpenStudy (anonymous):

What is 2.08 written as a fraction or mixed number?

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

im not to sure what to do

OpenStudy (anonymous):

@gottennis121

OpenStudy (gottennis121):

what's the whole number? (number before the decimal)

OpenStudy (anonymous):

2

OpenStudy (igreen):

0.08 is the same as 8/100, can you simplify that?

OpenStudy (anonymous):

ya

OpenStudy (igreen):

Okay, tell me what you get.

OpenStudy (anonymous):

wait wut i have to do?

OpenStudy (igreen):

Divide the top and bottom by 4. 8 / 4 = ? 100 / 4 = ?

OpenStudy (anonymous):

8 / 4 =2 100 / 4=25

OpenStudy (gottennis121):

http://www.mathsisfun.com/converting-decimals-fractions.html This might have additional help

OpenStudy (anonymous):

is meh answer D.?

OpenStudy (igreen):

Yes, so we have: 2 2/25

OpenStudy (igreen):

I don't think you gave choices.

OpenStudy (anonymous):

:D thx, but do i have to simplifiy

OpenStudy (anonymous):

oh i forgot to put the answer choices :/

OpenStudy (gottennis121):

That can not be simplified...

OpenStudy (anonymous):

okie

OpenStudy (igreen):

It's already simplified all the way.